Find the effective annual interest rate (EAR) for each of the following: 4.00 percent compounded quarterly. (Round answer to 2 decimal places, e.g. 15.25%.)

Effective annual rate %

4.80 percent compounded monthly. (Round answer to 2 decimal places, e.g. 15.25%.)

Effective annual rate%

6.50 percent compounded semiannually. (Round answer to 2 decimal places, e.g. 15.25%.)

Effective Annual Rate %

4.50 percent compounded daily. (Round answer to 2 decimal places, e.g. 15.25%. Use 365 days for calculation.)

Effective Annual Rate %
Answer #1

The effective annual rate is computed by using the below mentioned formula:

= (1 + r / m)m - 1

a. The effective annual rate is computed as follows:

= (1 + 0.04 / 4)4 - 1

= 4.06% Approximately

b. The effective annual rate is computed as follows:

= (1 + 0.048 / 12)12 - 1

= 4.91% Approximately

c. The effective annual rate is computed as follows:

= (1 + 0.065 / 2)2 - 1

= 6.61% Approximately

d. The effective annual rate is computed as follows:

= (1 + 0.045 / 365)365 - 1

= 4.60% Approximately

Feel free to ask in case of any query relating to this question